# EgoExo-Fitness: Towards Egocentric and Exocentric Full-Body Action Understanding

### v1

@@ 4.1 Action Classification
Train on Models Pretrain on Test on
Exo Ego
Exo 0.9194 0.0927
Ego I3D[9] K400[32] 0.1025 0.7469
Ego & Exo 0.8963 0.7266
Exo 0.9274 0.0836
Ego TSF[4] K600[8] 0.1417 0.7932
Ego & Exo 0.8894 0.7842
Exo 0.8940 0.0893
Ego EgoVLP[42] Ego4D[22] 0.0887 0.7977
Ego & Exo 0.8986 0.7932
Exo 0.8825 0.0814
Ego TSF[4] K600[8]+EgoExo4D[23] 0.1601 0.8000
Ego & Exo 0.8975 0.7840
Baseline Models. We apply three baseline models in Action Classification benchmark: 1) I3D [9] pretrained on K400 dataset [32]; 2) TimeSformer(TSF) [4] pretrained on K600 [8] and Ego-Exo4D [23] datasets; 3) EgoVLP [42] pretrained on Ego4D [22] dataset.
Experiment Results. Top-1 accuracies of different models are reported in Tab. 4. We analyze the results in the following aspects.
(1) Impacts of pretraining. Among all results, TimeSformer and I3D pretrained on Kinetics datasets achieve the best two performances (0.9274 and 0.9194) on exocentric videos. Similarly, TimeSformer pertrained on Ego-Exo4D performs best (0.8000) on egocentric videos, closely followed by the one pretrained on Ego4D (0.7977). Such results are attributed to view-related pre-training datasets (i.e., Kinetics are exocentric datasets; Ego4D and Ego-Exo4D consist of various egocentric videos).
(2) Analysis on view gap. Not surprisingly, models trained ego-only or exo-only data suffer from a significant performance drop on cross-view testing. Additionally, we find that mixing up cross-view data (Ego & Exo) for training does not always bring performance improvement. For I3D and TimeSformer pretrained on Kinetics datasets, performance drops on both egocentric and exocentric data. For TimeSformer pretrained from Ego4D and Ego-Exo4D, only performance on exocentric data obtains improvement when mixing up cross-view data for training. Such results indicate a great domain gap between ego-videos and exo-videos.
(3) Why do models perform worse on ego-videos? From Tab. 4, we also observe that models always perform worse on ego-videos than on exo-videos. We think this is because it is easier to observe similar action patterns from egocentric videos, which confuse models. Another reason is that it is more difficult to find discriminating clues from the Ego-M camera. Sec. A3.1 will show more analysis supporting these views.
@@ 4.2 Cross-view Sequence Verification
Sequence Verification (SV) [53, 14, 25] is proposed to verify the action order consistency of sequential videos under a scenario where precise temporal annotations are not provided, which shows great potential in video abstraction, industrial safety, and skill studying. Existing SV datasets [53] are collected either from exocentric cameras (i.e., COIN-SV and Diving48-SV) or from egocentric cameras (i.e., CSV), which constrains existing studies in an inner-view manner. However, in practical application, it is desirable to study whether a model can perform promising verification of two videos from egocentric and exocentric views. For instance, during our daily fitness exercises, an AI assistant in our eyewear can remind us whether we have missed any exercise program by verifying the sequence of exocentric expertise exemplar videos and the self-recorded egocentric videos. Motivated, we extend the traditional SV to Cross-View SV (CVSV).
@@
Task Setups. Cross-view Sequence verification (CVSV) aims to verify whether two fitness sequence videos have identical procedures. Two action sequence videos executing the same steps in the same order form a positive pair; otherwise, they are negative. The method should give a verification distance between each video pair and give the prediction by thresholding the distance.
@@
CVSV is more challenging than traditional SV because videos can be shot from either egocentric cameras or exocentric cameras. Hence, it is crucial for models to learn retrievable (or translatable) representations across views. More formal task setups will be introduced in Sec. A3.2.
Baseline model. We use the state-of-the-art SV model CAT [53] to conduct experiments. During training, CAT will take a pair of videos with the same action sequence as input. Except for classification loss, an extra sequence alignment loss is utilized to align the representations of each video in the pair.
@@
Metrics. (1) AUC. Following existing works [53, 14, 25], we first adopt the Area Under ROC Curve (AUC) to evaluate the performance. (2) Rank 1 & mAP. To further study the relationships among learned representations, we borrow the idea of image retrieval [7, 30, 34] and propose to use Rank-1 and mAP to evaluate CVSV models.
Experiment Results. Benchmark results are reported in Tab. 5 and Tab. 6. We analyze the results in the following aspects.
@@
(1) Influence of cross-view training data. As the first attempt, we wonder how cross-view training data will influence the performance. Hence, we separate all training video pairs into three parts (i.e., Exo-Exo, Ego-Ego and Exo-Ego) to study how cross-view training data would influence the performance. The results are shown in Tab. 5. First, we observe that combining all training pairs will benefit performance on exo-only and ego-exo pairs but bring a performance drop on ego-only pairs, which further indicates the domain gap between different views. Furthermore, compared with 0.8033 on ego-only data and 0.8221 on exo-only data, the best SV performance on ego-exo data is 0.7755, which indicates that cross-view sequence verification is a challenging task. Retrieval results also support this conclusion. Cross-view retrieval achieves much poorer performance (0.3 on Rank 1 and 0.25 on mAP) compared with inner-view retrieval.
(2) How many egocentric data is needed for CVSV? In practical application, it is much easier to collect exocentric videos than egocentric videos. Hence, it is desirable to study if a CVSV model can achieve superior performance with limited egocentric training videos. To the end, we gradually prune egocentric videos from the training set (i.e., 0%, 30%, 70% and 100%) and evaluate the performance. As shown in Tab. 6, when gradually prone training data of egocentric videos, the performance drops on all three metrics, which poses a great challenge for future study on settings with unbalanced (i.e., limited egocentric videos and rich exocentric videos) cross-view data.
@@
Prune rate AUC Rank 1 mAP
@@ 4.3 Guidance-based Execution Verification
Existing works in Action Assessment mainly focus on predicting the final score of an action video or a pair-wise ranking between a pair of videos. However, in real-world action guiding scenarios, providing interpretable feedback is more valuable than giving a score or a ranking. For example, our fitness coach will tell us which technical keypoints we are not satisfied with our execution, which will not only explain how well we have performed but also let us know how to improve. For the fitness coach, it is easy to compare the execution of an action with a technical keypoint and conclude whether the action satisfies it. However, such an ability has never been explored in action assessment. To address this issue, we make the first attempt to study interpretable action assessment and propose a novel task, Guidance-based Execution Verification (EVG).
Task Setups. We define guidance as a set of technical keypoints in text. Given the guidance, the goal of EVG is to verify whether the execution of an action satisfies the keypoints in the guidance. Formally, given an action video $v$ and $n$ technical keypoints $Q=\{q_{1},q_{2},...,q_{n}\}$, a model $F$ is asked to perform an $n$-way score prediction $P=\{p_{1},...,p_{n}\}=F(v,Q)$, where $p_{i}$ represents the verification score of the $i$-th keypoint. The higher the $p_{i}$, the more likely the action satisfies the $i$-th keypoint. During inference, a threshold $\tau$ is adopted to verify whether the action satisfies the keypoints. If $p_{i}$ > $\tau$, the model will predict the action “satisfies” the $i$-th keypoint. Otherwise, the model will return a result of being “unsatisfies”.
@@
In our implementation, visual and text encoders are implemented as the image and text encoders of pretrained CLIP [54] and frozen during training. The TCM module is designed as a Transformer Encoder and the CMM contains a Transformer Decoder together with a linear evaluator. For more details, please refer to the Sec. A3.3.
Table 7: Guidance-based Execution Verification results on different baselines. We report performance on egocentric and exocentric data alone with the average F1-score. “ego/exo” indicates independent models are trained on ego-only and exo-only data. “ego+exo” indicates that the model is trained on both ego- and exo- data. The full model of the proposed GEVFormer is colored in yellow.
Method Exo Ego Avg
@@
Experiment Results. We compare GEVFormer with four other naive methods and invariants: 1) Random: randomly predict an action satisfies a technical keypoint with 50% probability; 2) Distribution Prior: randomly predict an action meets a technical keypoint with the distribution prior; 3) CLIP-GEV: simply concat average-pooled visual embedding and text embeddings extracted by CLIP [54] and feed it into a linear evaluator to predict the results; 4) GEVFormer w/o alignment: ablate the alignment loss from GEVFormer. We find that in EgoExo-Fitness, samples with “satisfies” labels take up a much higher proportion than those with “unsatisfies” labels. We adopt the Precision, Recall, and F1-score to evaluate the model with the assumption that “unsatisfies” is the positive label.
As shown in Tab. 7, GEVFormer outperforms all naive baselines. Additionally, compared with the variants of GEVFormer, we have the same findings as on the other tasks that jointly training model on ego- and exo-centric data will not bring stable improvement (achieving 0.0087 improvement on egocentric data with 0.0192 drop on exocentric data). Surprisingly, when further adopting the alignment loss during training, GEVFormer can achieve the best performance on an egocentric dataset with 0.5425 F1-score, suffering only a 0.0022 performance drop on exocentric data.
@@ 5 Conclusion
We believe that studying egocentric full-body action understanding will benefit the development of AI-assistant. To enable this line of research, we focus on the scenario of fitness exercise and guiding, and introduce EgoExo-Fitness. With a diverse range of synchronized ego- and exo-centric fitness action sequence videos and rich annotations on temporal boundaries and interpretable action judgement, EgoExo-Fitness provides new resources for egocentric and exocentric full-body action understanding. To facilitate future research on EgoExo-Fitness, we construct benchmarks on five relevant tasks. Through experiment analysis, we evaluate the performance of baseline models and point out several interesting problems that await future research (e.g., how to better address cross-view modeling with unbalanced data; how to leverage synchronized exocentric data to achieve better performance on egocentric data).
@@
Recording System. For the two Go3 cameras, we use $2560\times 1440$ pixel resolution RGB images. For the GoPro camera, we use $1920\times 1080$ pixel resolution RGB images. For the side and front exocentric cameras, we set the resolution to be $1024\times 576$ and $1280\times 720$, respectively. After synchronization, video frames will be extracted with 30 FPS and resized to $456\times 256$.
Participants. We recruited 49 adults (35 males, 14 females) for data collection. Each participant was asked to participate at most nine rounds of recordings.
Action sequences. EgoExo-Fitness records 76 types of fitness action sequences, each containing 3 to 6 continuous fitness actions. Tab. A.1 provides the details of each action sequence.
@@
Privacy and Ethics. From the onset, privacy and ethics standards were critical to the data collection and release effort. All videos are recorded after we obtain the consent provided by participants. All human experts are asked to sign a privacy protection agreement to prevent data and privacy disclosure during the annotation process. To further protect the privacy and personal information, before the data release, we will ensure that the release resources do not contain privacy-sensitive content (e.g., real name, phone screens).
Table A.1: All recorded fitness sequences. For the correspondence between action names and abbreviations, please refer to Tab. 1. SID: sequence ID.
SIDs Action Orders SIDs Action Orders
  1 KPU, SU, JJ   39 LRL, SB, CJ
  2 KPU, JJ, SU   40 PU, LRL, KRAMC, CJ
  3 KTT, SU, JJ   41 PU, LRL, CJ, KRAMC
  4 KTT, JJ, SU   42 LRL, PU, KRAMC, CJ
  5 SU, KTT, JJ   43 SB, LRL, KRAMC, CJ
  6 KPU, JJ, HK   44 SB, LRL, CJ, KRAMC
  7 KPU, HK, JJ   45 KRAMC, LRL, SB, CJ
  8 JJ, KPU, HK   46 PU, SB, CJ, SS
  9 KPU, SU, LLKL, JJ   47 PU, SB, SS, CJ
  10 KPU, SU, JJ, LLKL   48 SB, PU, CJ, SS
  11 SU, KPU, LLKL, JJ   49 PU, SB, LRL, KRAMC, CJ
  12 KTT, SU, LLKL, JJ   50 PU, SB, LRL, CJ, KRAMC
  13 KTT, SU, JJ, LLKL   51 PU, LRL, SB, KRAMC, CJ
  14 LLKL, SU, KTT, JJ   52 SB, PU, LRL, CJ, SS
  15 KPU, KTT, JJ, HK   53 PU, SB, LRL, CJ, SS
  16 KPU, KTT, HK, JJ   54 PU, SB, SS, CJ, LRL
  17 KTT, KPU, JJ, HK   55 SB, HK, KRAMC, CJ, SS
  18 KPU, KTT, SU, LLKL, JJ   56 SB, KTT, KRAMC, CJ, SS
  19 KPU, KTT, SU, JJ, LLKL   57 SB, KPU, KRAMC, CJ, SS
  20 KPU, SU, KTT, LLKL, JJ   58 SB, CJ, KRAMC, KPU, SS
  21 KTT, KPU, SU, JJ, HK   59 SB, CJ, KRAMC, HK, SS
  22 KPU, KTT, SU, JJ, HK   60 SB, CJ, KRAMC, KTT, SS
  23 KPU, KTT, HK, JJ, SU   61 SB, CJ, KRAMC, LLKL, SS
  24 KTT, KRAMC, LLKL, JJ, HK   62 SB, CJ, KRAMC, SU, SS
  25 KTT, JJ, LLKL, KRAMC, HK   63 JJ, SB, KRAMC, CJ, SS
  26 KRAMC, KTT, LLKL, JJ, HK   64 LLKL, SB, KRAMC, CJ, SS
  27 KPU, KTT, SU, LLKL, JJ, HK   65 PU, SB, LRL, KRAMC, CJ, SS
  28 KPU, KTT, SU, JJ, HK, LLKL   66 PU, SB, LRL, CJ, SS, KRAMC
  29 KPU, SU, KTT, LLKL, JJ, HK   67 PU, LRL, SB, KRAMC, CJ, SS
  30 KTT, KPU, KRAMC, SU, JJ, HK   68 SB, PU, SU, LRL, CJ, SS
  31 KPU, KTT, KRAMC, SU, JJ, HK   69 SB, PU, LLKL, LRL, CJ, SS
  32 KPU, KTT, KRAMC, HK, JJ, SU   70 PU, SB, SU, SS, CJ, LRL
  33 KTT, KRAMC, SU, LLKL, JJ, HK   71 PU, SB, JJ, SS, CJ, LRL
  34 KTT, JJ, SU, LLKL, KRAMC, HK   72 PU, SB, LLKL, SS, CJ, LRL
  35 KRAMC, KTT, SU, LLKL, JJ, HK   73 PU, SB, KTT, SS, CJ, LRL
  36 PU, LRL, CJ   74 SB, HK, LRL, KRAMC, CJ, SS
  37 LRL, PU, CJ   75 SB, CJ, LRL, KRAMC, JJ, SS
  38 SB, CJ, LRL   76 SU, SB, LRL, KRAMC, CJ, SS

### v2

@@ 4.1 Action Classification
Baseline Models. We apply three baseline models in Action Classification benchmark: (1) I3D [9] pre-trained on K400 dataset [33]; (2) TimeSformer(TSF) [4] pre-trained on K600 [8] and Ego-Exo4D(EE4D) [24] datasets; (3) EgoVLP [43] pre-trained on Ego4D(E4D) [23] dataset.
@@
Train on Models Pretrain Test on Train on Models Pretrain Test on
Exo Ego Exo Ego
Exo 0.9194 0.0927 Exo 0.8940 0.0893
Ego I3D[9] K400[33] 0.1025 0.7469 Ego EgoVLP[43] Ego4D[23] 0.0887 0.7977
Ego & Exo 0.8963 0.7266 Ego & Exo 0.8986 0.7932
Exo 0.9274 0.0836 Exo K600[8] 0.8825 0.0814
Ego TSF[4] K600[8] 0.1417 0.7932 Ego TSF[4] + 0.1601 0.8000
Ego & Exo 0.8894 0.7842 Ego & Exo Ego-Exo4D[24] 0.8975 0.7840
Experiment Results. Top-1 accuracies of different models are reported in Tab. 4. We analyze the results in the following aspects. (1) Impacts of pretraining. Among all results, TSF and I3D pre-trained on Kinetics datasets achieve the best two performances (0.9274 and 0.9194) on exocentric videos. Similarly, TSF pre-trained on EE4D performs best (0.8000) on egocentric videos, closely followed by the one pre-trained on E4D (0.7977). Such results are attributed to view-related pre-training datasets (i.e., Kinetics are exocentric datasets; E4D and EE4D consist of various egocentric videos). (2) Analysis on the view gap. Not surprisingly, models trained ego-only or exo-only data suffer from a significant performance drop on cross-view testing. Additionally, we find that mixing up cross-view data (Ego & Exo) for training does not always bring performance improvement. For I3D and TSF pre-trained on Kinetics datasets, performance drops on both egocentric and exocentric data. For TSF pre-trained from E4D and EE4D, only performance on exocentric data obtains improvement when mixing up cross-view data for training. Such results indicate a great domain gap between ego-videos and exo-videos. (3) Why do models perform worse on ego-videos? From Tab. 4, we also observe that models always perform worse on ego-videos than on exo-videos. We think this is because it is easier to observe similar action patterns from egocentric videos, which confuse models. Another reason is that it is more difficult to find discriminating clues from the Ego-M camera. Appendix A3.1 provides more analysis supporting these views.
@@ 4.2 Cross-view Sequence Verification
Sequence Verification (SV) [55, 14, 26] is proposed to verify the action order consistency of sequential videos under a scenario where precise temporal annotations are not provided, which shows great potential in video abstraction, industrial safety, and skill studying. Existing SV datasets [55] are collected either from exocentric or egocentric cameras, which constrains existing studies in an inner-view manner. However, it is desirable to study whether a model can perform promising verification of two videos from egocentric and exocentric views. For instance, during our daily fitness exercises, an AI assistant in our eyewear can remind us whether we have missed any exercise program by verifying the sequence of exocentric expertise exemplar videos and the self-recorded egocentric videos. Hence, we extend the traditional SV to Cross-View SV (CVSV).
@@
Task Setups. CVSV aims to verify whether two fitness sequence videos have identical procedures. Two action sequence videos executing the same steps in the same order form a positive pair; otherwise, they are negative. The method should give a verification distance between each video pair based on the video representations, and give the prediction by thresholding the distance.
@@
CVSV is more challenging than traditional SV because videos can be shot from either egocentric or exocentric cameras. Hence, it is crucial for models to learn retrievable (or translatable) representations across views. More formal task setups will be introduced in Appendix A3.2.
Baseline model. We use the state-of-the-art SV model CAT [55] to conduct experiments. More details about the CAT will be introduced in Appendix A3.2.
@@
Metrics. (1) AUC: Following existing works [55, 14, 26], we first adopt the Area Under ROC Curve to evaluate the performance. (2) Rank 1 & mAP: To further study the relations among learned representations, we borrow the idea of image retrieval [7, 31, 35] and use Rank-1 and mAP to evaluate CVSV models.
@@
Experiment Results. Benchmark results are reported in Tab. 5 and Tab. 6. We analyze the results in the following aspects. (1) Influence of cross-view training data. As the first attempt, we wonder how cross-view training data will influence the performance. Hence, we separate all training video pairs into three parts (i.e., Exo-Exo, Ego-Ego and Exo-Ego) to study how cross-view training data would influence the performance. The results are shown in Tab. 5. First, we observe that combining all training pairs will benefit performance on exo-only and ego-exo pairs but bring a performance drop on ego-only pairs, which further indicates the domain gap between different views. Furthermore, compared with 0.8033 on ego-only data and 0.8221 on exo-only data, the best SV performance on ego-exo data is 0.7755, which indicates that cross-view sequence verification is a challenging task. Retrieval results also support this conclusion. Cross-view retrieval achieves much poorer performance (0.3 on Rank 1 and 0.25 on mAP) compared with inner-view retrieval. (2) How many egocentric data is needed for CVSV? In practical application, it is much easier to collect exocentric videos than egocentric videos. Hence, it is desirable to study if a CVSV model can achieve superior performance with limited egocentric training videos. To the end, we gradually prune egocentric videos from the training set (i.e., 0%, 30%, 70% and 100%) and evaluate the performance. As shown in Tab. 6, when gradually prone training data of egocentric videos, the performance drops on all metrics, posing a great challenge for future study on settings with unbalanced (i.e., limited egocentric videos and rich exocentric videos) cross-view data.
@@
Prune AUC Rank 1 mAP Prune AUC Rank 1 mAP
@@ 4.3 Guidance-based Execution Verification
Existing works in Action Assessment mainly focus on predicting the final score of an action video or a pair-wise ranking between a pair of videos. However, in real-world action guiding scenarios, providing interpretable feedback is more valuable than giving a score or a ranking. For example, our fitness coach will tell us which technical keypoints are not satisfied in our executions, which will not only explain how well we have performed but also let us know how to improve. However, such an ability has never been explored in action assessment. To address this issue, we make the first attempt to study interpretable action assessment and propose a novel task termed Guidance-based Execution Verification (GEV).
Task Setups. Given a set of technical keypoints in text as the guidance, the goal of GEV is to verify whether the execution of an action satisfies the keypoints in the guidance. Formally, given an action video $v$ and $n$ technical keypoints $Q=\{q_{1},q_{2},...,q_{n}\}$, a model $F$ is asked to perform an $n$-way score prediction $P=\{p_{1},...,p_{n}\}=F(v,Q)$, where $p_{i}$ represents the verification score of the $i$-th keypoint. The higher the $p_{i}$, the more likely the action satisfies the $i$-th keypoint. During inference, a threshold $\tau$ is adopted to verify whether the action satisfies the keypoints. If $p_{i}$ > $\tau$, the model predicts the action “satisfies” the $i$-th keypoint. Otherwise, the model returns a result of being “unsatisfies”.
@@
In our implementation, visual and text encoders are designed as the image and text encoders of pre-trained CLIP [56] and frozen during training. The TCM is designed as a Transformer Encoder and the CMV contains a Transformer Decoder together with a linear evaluator. See more details in Appendix A3.3.
Table 7: Guidance-based execution verification results on different baselines. We report performance on ego- and exo-centric data alone with the average F1-score. “ego/exo” indicates ndependent models are trained on ego-only and exo-only data. “ego+exo” indicates the model is trained on both ego-and-exo data.
Methods Exo Ego Avg
@@
Experiment Results. We compare GEVFormer with four other naive methods and invariants: (1) Random: Randomly predict an action satisfies a technical keypoint with 50% probability; (2) Distribution Prior: Randomly predict an action satisfies a keypoint with the distribution prior; (3) CLIP-GEV: Simply concatenate average-pooled visual embedding and text embeddings extracted by CLIP [56] and feed it into a linear evaluator to predict the results; (4) GEVFormer w/o alignment: Ablate $L_{Align}$ from GEVFormer. To evaluate the methods, we adopt the Precision, Recall, and F1-score. Here we regard “unsatisfies” as the positive label since samples with “satisfies” labels take up a much higher proportion than those with “unsatisfies” labels in EgoExo-Fitness.
As shown in Tab. 7, GEVFormer outperforms all naive baselines. Besides, compared with the variants of GEVFormer, we have the same findings as on the other tasks that jointly training models on ego- and exo-centric data will not bring stable improvement (achieving 0.0087 improvement on egocxentric data with 0.0192 drop on exocentric data). Surprisingly, when further adopting the $L_{Align}$, GEVFormer achieves the best performance on egocentric data with 0.5425 F1-score, suffering only a 0.0022 performance drop on exocentric data.
@@ 5 Conclusion
We believe that studying egocentric full-body action understanding will benefit the development of AI-assistant. To enable this line of research, we focus on the scenario of fitness exercise and guiding, and introduce EgoExo-Fitness. With a diverse range of synchronized ego- and exo-centric fitness action sequence videos and rich annotations on temporal boundaries and interpretable action judgement, EgoExo-Fitness provides new resources for egocentric and exocentric full-body action understanding. To facilitate future research on EgoExo-Fitness, we construct benchmarks on five relevant tasks. Through experiment analysis, we evaluate the performance of baseline models and point out several interesting problems that await future research (e.g., how to better address cross-view modeling with unbalanced data; how to leverage synchronized exocentric data to achieve better performance).
Acknowledgments
This work was supported partially by the National Key Research and Development Program of China (2023YFA1008503), NSFC(U21A20471, 62206315), Guangdong NSF Project (No. 2023B1515040025, 2020B1515120085, 2024A15150-10101), Guangzhou Basic and Applied Basic Research Scheme(2024A04J4067). The authors thank Kun-Yu Lin for the valuable discussions. The authors also thank anonymous reviewers and ACs for their constructive suggestions.
@@ Acknowledgments
Recording System. For the two Insta-Go3 cameras, we use $2560\times 1440$ pixel resolution RGB images. For the GoPro camera, we use $1920\times 1080$ pixel resolution RGB images. For the side and front exocentric cameras, we set the resolution to be $1024\times 576$ and $1280\times 720$, respectively. After synchronization, video frames will be extracted with 30 FPS and resized to $456\times 256$.
Participants. We recruited 40 adults (28 males, 12 females) for data collection. Each participant was asked to participate at most nine rounds of recordings.
Action sequences. EgoExo-Fitness records 86 types of fitness action sequences, each containing 3 to 6 continuous fitness actions. Tab. A.2 provides the details of each action sequence.
@@
Privacy and Ethics. From the onset, privacy and ethics standards were critical to the data collection and release effort. All videos are recorded after we obtain the consent provided by participants. All human experts are asked to sign a privacy protection agreement to prevent data and privacy disclosure during the annotation process. To further protect the privacy and personal information, before the data release, we will ensure that the release resources do not contain privacy-sensitive content (e.g., real name).
Table A.1: Recorded fitness actions. Abbr.: the abbreviation of the fitness action.
Action types Abbr. Action types Abbr. Action types Abbr.
1: Kneeling Push-ups KPU 5: Shoulder Bridge SB 10: Jumping Jacks JJ
2: Push-ups PU 6: Sit-ups SU 11: High Knee HK
3: Kneeling Torso Twist KTT 7: Leg Reverse Lunge LRL 12: Clap Jacks CJ
4: Knee Raise and Abdomi- KRAMC 8: Leg Lunge with Knee Lift LLKL
         nal Muscles Contract 9: Sumo Squat SS
Table A.2: All recorded fitness sequences. For the correspondence between action names and abbreviations, please refer to Tab. A.1. SID: sequence ID.
SIDs Action Orders SIDs Action Orders SIDs Action Orders
1 KPU, SU, JJ 30 KTT, KPU, KRAMC, SU, JJ, HK 59 SB, CJ, KRAMC, HK, SS
2 KPU, JJ, SU 31 KPU, KTT, KRAMC, SU, JJ, HK 60 SB, CJ, KRAMC, KTT, SS
3 KTT, SU, JJ 32 KPU, KTT, KRAMC, HK, JJ, SU 61 SB, CJ, KRAMC, LLKL, SS
4 KTT, JJ, SU 33 KTT, KRAMC, SU, LLKL, JJ, HK 62 SB, CJ, KRAMC, SU, SS
5 SU, KTT, JJ 34 KTT, JJ, SU, LLKL, KRAMC, HK 63 JJ, SB, KRAMC, CJ, SS
6 KPU, JJ, HK 35 KRAMC, KTT, SU, LLKL, JJ, HK 64 LLKL, SB, KRAMC, CJ, SS
7 KPU, HK, JJ 36 PU, LRL, CJ 65 PU, SB, LRL, KRAMC, CJ, SS
8 JJ, KPU, HK 37 LRL, PU, CJ 66 PU, SB, LRL, CJ, SS, KRAMC
9 KPU, SU, LLKL, JJ 38 SB, CJ, LRL 67 PU, LRL, SB, KRAMC, CJ, SS
10 KPU, SU, JJ, LLKL 39 LRL, SB, CJ 68 SB, PU, SU, LRL, CJ, SS
11 SU, KPU, LLKL, JJ 40 PU, LRL, KRAMC, CJ 69 SB, PU, LLKL, LRL, CJ, SS
12 KTT, SU, LLKL, JJ 41 PU, LRL, CJ, KRAMC 70 PU, SB, SU, SS, CJ, LRL
13 KTT, SU, JJ, LLKL 42 LRL, PU, KRAMC, CJ 71 PU, SB, JJ, SS, CJ, LRL
14 LLKL, SU, KTT, JJ 43 SB, LRL, KRAMC, CJ 72 PU, SB, LLKL, SS, CJ, LRL
15 KPU, KTT, JJ, HK 44 SB, LRL, CJ, KRAMC 73 PU, SB, KTT, SS, CJ, LRL
16 KPU, KTT, HK, JJ 45 KRAMC, LRL, SB, CJ 74 SB, HK, LRL, KRAMC, CJ, SS
17 KTT, KPU, JJ, HK 46 PU, SB, CJ, SS 75 SB, CJ, LRL, KRAMC, JJ, SS
18 KPU, KTT, SU, LLKL, JJ 47 PU, SB, SS, CJ 76 SU, SB, LRL, KRAMC, CJ, SS
19 KPU, KTT, SU, JJ, LLKL 48 SB, PU, CJ, SS 77 LLKL,KRAMC,HK
20 KPU, SU, KTT, LLKL, JJ 49 PU, SB, LRL, KRAMC, CJ 78 SB,LRL,CJ
21 KTT, KPU, SU, JJ, HK 50 PU, SB, LRL, CJ, KRAMC 79 PU,SB,CJ,LRL
22 KPU, KTT, SU, JJ, HK 51 PU, LRL, SB, KRAMC, CJ 80 KTT,SU,JJ,HK
23 KPU, KTT, HK, JJ, SU 52 SB, PU, LRL, CJ, SS 81 KPU,KTT,SU,JJ
24 KTT, KRAMC, LLKL, JJ, HK 53 PU, SB, LRL, CJ, SS 82 KPU,KTT,LLKL,SU,JJ,HK
25 KTT, JJ, LLKL, KRAMC, HK 54 PU, SB, SS, CJ, LRL 83 KPU,KTT,JJ,HK,LLKL
26 KRAMC, KTT, LLKL, JJ, HK 55 SB, HK, KRAMC, CJ, SS 84 PU,SB,KRAMC,CJ,SS
27 KPU, KTT, SU, LLKL, JJ, HK 56 SB, KTT, KRAMC, CJ, SS 85 KPU,SU,KTT,KRAMC,JJ
28 KPU, KTT, SU, JJ, HK, LLKL 57 SB, KPU, KRAMC, CJ, SS 86 KTT,KRAMC,LLKL,SU,JJ,HK
29 KPU, SU, KTT, LLKL, JJ, HK 58 SB, CJ, KRAMC, KPU, SS